# Universidad Andres Bello, Facultad de Ciencias Exactas, Departamento de Ciencias Físicas - Instituto de Astrofísica, Autopista Concepcion-Talcahuano 7100, Talcahuano, Chile Instituto de Física, Universidad de Antioquia, Calle 70 52-21, Medellín, Colombia

### v1

@@ 5 Conclusions
We obtained a mean metallicity of [Fe/H]= 0.87 0.02 for the cluster, with no clear evidence of any intrinsic Fe abundance variation, in very good agreement with the results of Deras et al. (2023). This value is 0.5 dex more metal poor than the values traditionally derived for this cluster (Vásquez et al., 2018; Harris, 1996, 2010 edition). We encourage the use of this new derived metallicity above the previous ones since our determination involves a large sample of stars observed at high resolution and SNR. Also, since our metallicity supports that derived by Deras, the age derived by them (13.1 ± 0.5 Gyrs) is therefore supported by this study.
@@
We provide abundances for C, N, O, Mg, Al, Si, P, K, Ca, Ti, V, Cr, Mn, Ni and Ce. This is the first non-automatic spectroscopic abundance determination for this cluster, hence the most reliable chemical study to date for this cluster, preferable to previous studies like Schiavon et al. (2024) and Conroy et al. (2018), who used ASPCAP and stellar population models, respectively. With the abundances of our available -elements we derived [/Fe] = 0.3 0.02.
@@
The C and N levels of the stars show a clear anticorrelation, with a noticeable enhancement in Nitrogen, indicating the presence of MPs within the cluster.
@@
The isochrone fitting included in our method for determining the atmospheric parameters provided the following information for NGC 6316: AV = 1.91 0.05, RV = 2.7 0.1 (hence giving a higher E(B-V) = 0.71 than other studies that use RV = 3.1), (m-M)0 = 15.32 0.05 and [M/H] = 0.6 0.05. This latter is in good agreement with our global metallicity derived from our [Fe/H] and [/Fe] abundances ([M/H] = 0.66 0.03) using the formula from Salaris, Chieffi, & Straniero (1993).
@@
A comparison of the abundances of NGC 6316 with those of other in-situ GCs shows a very good agreement in almost all cases, with NGC 6316 having higher spread in most cases except Al and Si.
@@
Although it is known that ASPCAP does not perform well for 2P stars, inflating the precision of mean abundances, the elements Al and K, and the iron-peak elements Ti, Cr, Mn and Ni show excellent agreement in the mean abundances and dispersion for our GC, while O and Mg show a good agreement only in their mean abundances. Vanadium is the only element that presents higher dispersion in ASPCAP abundances than in BACCHUS, while all the other elements present lower (C, N, O, Mg, Ca, Fe and Ce) or equal (Al, Si, K, Ti, Cr, Mn and Ni) dispersions than BACCHUS abundances.
7.
Replicating the scheme proposed by Belokurov & Kravtsov (2024), but based on the mean [Mg/Fe] and [Al/Fe] abundances of 1P stars only, NGC 6316 is classified as an in-situ GC. This study discards the reliability of using 2P stars. However, the separation line defined by BK24 was derived from both 1P and 2P stars. The authors acknowledge that a separation line derived exclusively from 1P stars should be implemented to ensure a more consistent and accurate analysis.
@@ Acknowledgements.
SV gratefully acknowledges the support provided by Fondecyt regular no. 1220264 and by the ANID BASAL projects FB210003. D.G. gratefully acknowledges the support provided by Fondecyt regular n. 1220264. D.G. also acknowledges financial support from the Dirección de Investigación y Desarrollo de la Universidad de La Serena through the Programa de Incentivo a la Investigación de Académicos (PIA-DIDULS). The team acknowledges J.G. Fernandez-Trincado for his help and commentaries. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.References
@@
- Allende Prieto et al. (2008) Allende Prieto C., Majewski S. R., Schiavon R., Cunha K., Frinchaboy P., Holtzman J., Johnston K., et al., 2008, AN, 329, 1018. doi:10.1002/asna.200811080
- Alonso et al. (1999) Alonso, A., Arribas, S., & Martínez-Roger, C. 1999, A&AS, 140, 261
@@
- Bajkova & Bobylev (2020) Bajkova A. T., Bobylev V. V., 2020, arXiv, arXiv:2008.13624
- Baumgardt et al. (2019) Baumgardt, H., Hilker, M., Sollima, A., et al. 2019, MNRAS, 482, 5138. doi:10.1093/mnras/sty2997
- Bastian et al. (2013) Bastian, N., Lamers, H. J. G. L. M., de Mink, S. E., et al. 2013, MNRAS, 436, 2398
- Bastian & Lardo (2018) Bastian N., Lardo C., 2018, ARA&A, 56, 83. doi:10.1146/annurev-astro-081817-051839
- Bedin et al. (2004) Bedin, L. R., Piotto, G., Anderson, J., et al. 2004, ApJ, 605, L125
- Bellazzini, Ferraro, & Ibata (2003) Bellazzini M., Ferraro F. R., Ibata R., 2003, AJ, 125, 188. doi:10.1086/344072
@@
- Bonatto, Campos, & Kepler (2013) Bonatto C., Campos F., Kepler S. O., 2013, MNRAS, 435, 263. doi:10.1093/mnras/stt1304
- Bovy et al. (2019) Bovy, J., Leung, H. W., Hunt, J. A. S., et al. 2019, MNRAS, 490, 4740. doi:10.1093/mnras/stz2891
@@
- Campbell et al. (2012) Campbell S. W., Yong D., Wylie-de Boer E. C., Stancliffe R. J., Lattanzio J. C., Angelou G. C., D’Orazi V., et al., 2012, ApJL, 761, L2. doi:10.1088/2041-8205/761/1/L2
- Cannon & Stobie (1973) Cannon R. D., Stobie R. S., 1973, MNRAS, 162, 207. doi:10.1093/mnras/162.3.207
- Canterna (1976) Canterna, R. 1976, AJ, 81, 228
@@
- Carretta et al. (2009) Carretta E., Bragaglia A., Gratton R. G., Lucatello S., Catanzaro G., Leone F., Bellazzini M., et al., 2009, A&A, 505, 117. doi:10.1051/0004-6361/200912096
- Carretta et al. (2009) Carretta E., Bragaglia A., Gratton R., Lucatello S., 2009, A&A, 505, 139. doi:10.1051/0004-6361/200912097
- Carretta et al. (2010) Carretta E., Bragaglia A., Gratton R. G., Recio-Blanco A., Lucatello S., D’Orazi V., Cassisi S., 2010, A&A, 516, A55. doi:10.1051/0004-6361/200913451
@@
- Cohen (2004) Cohen J. G., 2004, AJ, 127, 1545. doi:10.1086/382104
@@
- D’Ercole et al. (2008) D’Ercole, A., Vesperini, E., D’Antona, F., McMillan, S. L. W., & Recchi, S. 2008, MNRAS, 391, 825
- Dalessandro et al. (2018) Dalessandro E., Lardo C., Cadelano M., Saracino S., Bastian N., Mucciarelli A., Salaris M., et al., 2018, A&A, 618, A131. doi:10.1051/0004-6361/201833650
- Decressin et al. (2007) Decressin, T., Charbonnel, C., & Meynet, G. 2007, A&A, 475, 859
@@
- Dotter et al. (2011) Dotter, A., Sarajedini, A., & Anderson, J. 2011, ApJ, 738, 74
- Dotter et al. (2018) Dotter, A., Milone, A. P., Conroy, C., et al. 2018, ApJ, 865, L10
- Elmegreen (2017) Elmegreen B. G., 2017, ApJ, 836, 80. doi:10.3847/1538-4357/836/1/80
- Fernie (1983) Fernie, J. D. 1983, PASP, 95, 782
@@
- Frelijj et al. (2017) Frelijj H., Geisler D., Cummings J., Cohen R. E., Mauro F., Munoz C., Villanova S., et al., 2017, MNRAS, 472, 4532. doi:10.1093/mnras/stx2280
@@
- Frelijj et al. (2022) Frelijj H., Geisler D., Villanova S., Munoz C., 2022, RMxAA, 58, 181. doi:10.22201/ia.01851101p.2022.58.02.02
@@
- Geisler, Claria, & Minniti (1991) Geisler D., Claria J. J., Minniti D., 1991, AJ, 102, 1836. doi:10.1086/116008
- Geisler (1996) Geisler D., 1996, AJ, 111, 480. doi:10.1086/117799
@@
- Gieles et al. (2018) Gieles M., Charbonnel C., Krause M. G. H., Hénault-Brunet V., Agertz O., Lamers H. J. G. L. M., Bastian N., et al., 2018, MNRAS, 478, 2461. doi:10.1093/mnras/sty1059
@@
- Gratton, Carretta, & Castelli (1996) Gratton R. G., Carretta E., Castelli F., 1996, A&A, 314, 191. doi:10.48550/arXiv.astro-ph/9603011
@@
- Hamuy et al. (2006) Hamuy M., Folatelli G., Morrell N. I., Phillips M. M., Suntzeff N. B., Persson S. E., Roth M., et al., 2006, PASP, 118, 2. doi:10.1086/500228
- Han et al. (2009) Han S.-I., Lee Y.-W., Joo S.-J., Sohn S. T., Yoon S.-J., Kim H.-S., Lee J.-W., 2009, ApJL, 707, L190. doi:10.1088/0004-637X/707/2/L190
@@
- Johnson et al. (2017) Johnson C. I., Caldwell N., Rich R. M., Mateo M., Bailey J. I., Clarkson W. I., Olszewski E. W., et al., 2017, ApJ, 836, 168. doi:10.3847/1538-4357/836/2/168.
@@
- Koch & Caffau (2011) Koch A., Caffau E., 2011, A&A, 534, A52. doi:10.1051/0004-6361/201117678
- Kurucz (1970) Kurucz, R. L. 1970, SAO Special Report, 309
- Law & Majewski (2010) Law D. R., Majewski S. R., 2010, ApJ, 718, 1128. doi:10.1088/0004-637X/718/2/1128
@@
- Milone et al. (2009) Milone A. P., Stetson P. B., Piotto G., Bedin L. R., Anderson J., Cassisi S., Salaris M., 2009, A&A, 503, 755. doi:10.1051/0004-6361/200912256
@@
- Monaco et al. (2018) Monaco L., Villanova S., Carraro G., Mucciarelli A., Moni Bidin C., 2018, A&A, 616, A181. doi:10.1051/0004-6361/201832979
- Moreno, Pichardo, & Schuster (2015) Moreno E., Pichardo B., Schuster W. J., 2015, MNRAS, 451, 705. doi:10.1093/mnras/stv962
@@
- Muñoz et al. (2021) Muñoz C., Geisler D., Villanova S., Sarajedini A., Frelijj H., Vargas C., Monaco L., et al., 2021, MNRAS, 506, 4676. doi:10.1093/mnras/stab1891
@@
- O’Donnell (1994) O’Donnell J. E., 1994, ApJ, 422, 158. doi:10.1086/173713
- Parmentier & Pasquali (2022) Parmentier G., Pasquali A., 2022, ApJ, 924, 81. doi:10.3847/1538-4357/ac32d8
@@
- Ramírez & Meléndez (2005) Ramírez, I. & Meléndez, J. 2005, ApJ, 626, 465
- Renzini et al. (2015) Renzini A., D’Antona F., Cassisi S., King I. R., Milone A. P., Ventura P., Anderson J., et al., 2015, MNRAS, 454, 4197. doi:10.1093/mnras/stv2268
- Robin et al. (2003) Robin A. C., Reylé C., Derrière S., Picaud S., 2003, A&A, 409, 523. doi:10.1051/0004-6361:20031117
- Robin et al. (2014) Robin A. C., Reylé C., Fliri J., Czekaj M., Robert C. P., Martins A. M. M., 2014, A&A, 569, A13. doi:10.1051/0004-6361/201423415
@@
- Salinas, & Strader (2015) Salinas, R., & Strader, J. 2015, ApJ, 809, 169
- Sandage (1953) Sandage A. R., 1953, AJ, 58, 61. doi:10.1086/106822
- Sanders, Smith, & Evans (2019) Sanders J. L., Smith L., Evans N. W., 2019, MNRAS, 488, 4552. doi:10.1093/mnras/stz1827
@@
- Sbordone et al. (2005) Sbordone L., Bonifacio P., Marconi G., Buonanno R., Zaggia S., 2005, A&A, 437, 905. doi:10.1051/0004-6361:20042315
- Sbordone et al. (2007) Sbordone L., Bonifacio P., Buonanno R., Marconi G., Monaco L., Zaggia S., 2007, A&A, 465, 815. doi:10.1051/0004-6361:20066385
- Sbordone et al. (2011) Sbordone, L., Salaris, M., Weiss, A., & Cassisi, S. 2011, A&A, 534, A9
- Schiavon et al. (2017) Schiavon R. P., Johnson J. A., Frinchaboy P. M., Zasowski G., Mészáros S., García-Hernández D. A., Cohen R. E., et al., 2017, MNRAS, 466, 1010. doi:10.1093/mnras/stw3093
@@
- Sirianni et al. (2005) Sirianni, M., Jee, M. J., Benítez, N., et al. 2005, PASP, 117, 1049. doi:10.1086/444553
- Smith et al. (2021) Smith V. V., Bizyaev D., Cunha K., Shetrone M. D., Souto D., Allende Prieto C., Masseron T., et al., 2021, AJ, 161, 254. doi:10.3847/1538-3881/abefdc
- Smith et al. (2002) Smith, G. H., Sneden, C., & Kraft, R. P. 2002, AJ, 123, 1502. doi:10.1086/338855
- Sneden (1973) Sneden, C. 1973, ApJ, 184, 839. doi:10.1086/152374
- Stetson (1987) Stetson P. B., 1987, PASP, 99, 191. doi:10.1086/131977
- Stetson (1994) Stetson P. B., 1994, PASP, 106, 250. doi:10.1086/133378
- Taylor (2005) Taylor, M. B. 2005, Astronomical Data Analysis Software and Systems XIV, 347, 29
- Tautvaišienė et al. (2004) Tautvaišienė, G., Wallerstein, G., Geisler, D., et al. 2004, AJ, 127, 373
@@
- Walker et al. (2011) Walker A. R., Kunder A. M., Andreuzzi G., di Cecco A., Stetson P. B., Monelli M., Cassisi S., et al., 2011, MNRAS, 415, 643. doi:10.1111/j.1365-2966.2011.18736.x

### v2

@@ 5 Conclusions
We obtained a mean metallicity of [Fe/H]= 0.87 0.02 for the cluster, with no clear evidence of any intrinsic Fe abundance variation, in very good agreement with the results of D23. This value is 0.5 dex more metal poor than the values traditionally derived for this cluster (Vásquez et al., 2018, and H10). We encourage the use of this new derived metallicity above the previous ones since our determination involves a large sample of stars observed at high resolution and SNR. Also, since our metallicity supports that derived by D23, the age derived by them (13.1 ± 0.5 Gyrs) is therefore supported by this study.
@@
We provide abundances for C, N, O, Mg, Al, Si, P, K, Ca, Ti, V, Cr, Mn, Ni and Ce. This is the first non-automatic high resolution spectroscopic abundance determination for this cluster, hence the most reliable chemical study to date for this cluster, preferable to previous studies like Conroy et al. (2018), who used stellar population models, respectively. Using the abundances of Si and Ca -elements we derived [/Fe] = 0.31 0.02.
@@
The C and N levels of the stars show a clear anti-correlation, with a noticeable enhancement in Nitrogen, indicating the presence of MPs within the cluster.
@@
The isochrone fitting included in our method for determining the atmospheric parameters provided the following information for NGC 6316: AV = 1.91 0.05, RV = 2.7 0.1 (hence giving a higher E(B-V) = 0.71 than other studies that use RV = 3.1), (m-M)0 = 15.32 0.05 and [M/H] = 0.6 0.05. This latter is in good agreement with our global metallicity derived from our [Fe/H] and [/Fe] abundances ([M/H] = 0.65 0.08), using the formula from Salaris, Chieffi, & Straniero (1993).
@@
A comparison of the abundances of NGC 6316 with those of other in-situ GCs shows a very good agreement in almost all cases, with NGC 6316 having a higher spread in C, N and Ce.
@@
We found that the method proposed by BK24 to classify GCs based on their [Mg/Fe] vs. [Al/Fe] abundances did not function as expected. However, a classification using only 1P stars could work.
@@ Acknowledgements.
The authors acknowledge the referee’s comments that substantially improved the content of this paper. H.F. gratefully acknowledges the support provided by ALMA-ANID funds, No. 31230013. D.G.D. gratefully acknowledges the support provided by ALMA-ANID funds, No. 31230017. S.V. gratefully acknowledges the support provided by Fondecyt regular n. 1220264 and by the ANID BASAL projects FB210003. D.G. gratefully acknowledges the support provided by Fondecyt regular n. 1220264. D.G. also acknowledges financial support from the Dirección de Investigación y Desarrollo de la Universidad de La Serena through the Programa de Incentivo a la Investigación de Académicos (PIA-DIDULS). C.M. thanks the support provided by ANID GEMINI Postdoctorado No. 32230017. The team acknowledges J.G. Fernandez-Trincado for his help and commentaries. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.References
@@
- Armandroff & Zinn (1988) Armandroff T. E., Zinn R., 1988, AJ, 96, 92. doi:10.1086/114792
- Asplund, Grevesse, & Sauval (2005) Asplund M., Grevesse N., Sauval A. J., 2005, ASPC, 336, 25
- Barrera et al. (2025) Barrera N., Villanova S., Geisler D., 2025, A&A, accepted.
@@
- Caloi & D’Antona (2008) Caloi V., D’Antona F., 2008, ApJ, 673, 847. doi:10.1086/523346
@@
- Carretta et al. (2009a) Carretta E., Bragaglia A., Gratton R. G., Lucatello S., Catanzaro G., Leone F., Bellazzini M., et al., 2009, A&A, 505, 117. doi:10.1051/0004-6361/200912096
- Carretta et al. (2009b) Carretta E., Bragaglia A., Gratton R., Lucatello S., 2009, A&A, 505, 139. doi:10.1051/0004-6361/200912097
- Carretta et al. (2009c) Carretta E., Bragaglia A., Gratton R., D’Orazi V., Lucatello S., 2009, A&A, 508, 695. doi:10.1051/0004-6361/200913003
@@
- Cohen & Sarajedini (2012) Cohen R. E., Sarajedini A., 2012, MNRAS, 419, 342. doi:10.1111/j.1365-2966.2011.19697.x
@@
- De Silva et al. (2015) De Silva G. M., Freeman K. C., Bland-Hawthorn J., Martell S., de Boer E. W., Asplund M., Keller S., et al., 2015, MNRAS, 449, 2604. doi:10.1093/mnras/stv327
- Dias et al. (2016) Dias B., Barbuy B., Saviane I., Held E. V., Da Costa G. S., Ortolani S., Gullieuszik M., et al., 2016, A&A, 590, A9. doi:10.1051/0004-6361/201526765
- Dotter et al. (2010) Dotter A., Sarajedini A., Anderson J., Aparicio A., Bedin L. R., Chaboyer B., Majewski S., et al., 2010, ApJ, 708, 698. doi:10.1088/0004-637X/708/1/698
@@
- Geisler et al. (2025) Geisler D., Munoz C., Villanova S., et al., A&A, submitted.
- Goldsbury, Heyl, & Richer (2013) Goldsbury R., Heyl J., Richer H., 2013, ApJ, 778, 57. doi:10.1088/0004-637X/778/1/57
- Gonzalez et al. (2012) Gonzalez O. A., Rejkuba M., Zoccali M., Valenti E., Minniti D., Schultheis M., Tobar R., et al., 2012, A&A, 543, A13. doi:10.1051/0004-6361/201219222
@@
- Henao et al. (2025) Henao L., Villanova S., Geisler D., Fernández-Trincado J. G., 2025, A&A, 696, A154. doi:10.1051/0004-6361/202451793
@@
- Momany et al. (2012) Momany Y., Saviane I., Smette A., Bayo A., Girardi L., Marconi G., Milone A. P., et al., 2012, A&A, 537, A2. doi:10.1051/0004-6361/201117223
- Muñoz et al. (2017) Muñoz C., Villanova S., Geisler D., Saviane I., Dias B., Cohen R. E., Mauro F., 2017, A&A, 605, A12. doi:10.1051/0004-6361/201730468
@@
- Nataf et al. (2013) Nataf D. M., Gould A., Fouqué P., Gonzalez O. A., Johnson J. A., Skowron J., Udalski A., et al., 2013, ApJ, 769, 88. doi:10.1088/0004-637X/769/2/88
@@
- Origlia et al. (2005) Origlia L., Valenti E., Rich R. M., Ferraro F. R., 2005, MNRAS, 363, 897. doi:10.1111/j.1365-2966.2005.09490.x
- Paust et al. (2010) Paust N. E. Q., Reid I. N., Piotto G., Aparicio A., Anderson J., Sarajedini A., Bedin L. R., et al., 2010, AJ, 139, 476. doi:10.1088/0004-6256/139/2/476
@@
- Reed, Hesser, & Shawl (1988) Reed B. C., Hesser J. E., Shawl S. J., 1988, PASP, 100, 545. doi:10.1086/132202
@@
- Sarajedini et al. (2007) Sarajedini A., Bedin L. R., Chaboyer B., Dotter A., Siegel M., Anderson J., Aparicio A., et al., 2007, AJ, 133, 1658. doi:10.1086/511979
- Saviane et al. (2012) Saviane I., Da Costa G. S., Held E. V., Sommariva V., Gullieuszik M., Barbuy B., Ortolani S., 2012, A&A, 540, A27. doi:10.1051/0004-6361/201118138
@@
- Valenti, Ferraro, & Origlia (2007) Valenti E., Ferraro F. R., Origlia L., 2007, AJ, 133, 1287. doi:10.1086/511271
- Valenti, Origlia, & Rich (2011) Valenti E., Origlia L., Rich R. M., 2011, MNRAS, 414, 2690. doi:10.1111/j.1365-2966.2011.18580.x
- VandenBerg et al. (2013) VandenBerg D. A., Brogaard K., Leaman R., Casagrande L., 2013, ApJ, 775, 134. doi:10.1088/0004-637X/775/2/134
@@
- Vásquez et al. (2015) Vásquez S., Zoccali M., Hill V., Gonzalez O. A., Saviane I., Rejkuba M., Battaglia G., 2015, A&A, 580, A121. doi:10.1051/0004-6361/201526534
@@
- Villanova et al. (2017) Villanova S., Moni Bidin C., Mauro F., Munoz C., Monaco L., 2017, MNRAS, 464, 2730. doi:10.1093/mnras/stw2509
@@
- Zinn & West (1984) Zinn R., West M. J., 1984, ApJS, 55, 45. doi:10.1086/190947
- Zinn (1985) Zinn R., 1985, ApJ, 293, 424. doi:10.1086/163249
Appendix A Comparisons with ASPCAP
Given the substantial number of new abundances obtained, it is a good opportunity to compare them with the ASPCAP values provided in the APOGEE value-added catalogue from Schiavon et al. (2024, hereafter S24). A similar analysis, expanded for all CAPOS cluster, is being carried out in Geisler et al. (2025). S24 presented abundances for a group of stars in NGC 6316, which includes all 10 of our target stars, enabling a reliable comparison. For brevity, we will refer to our results as BACCHUS and those from S24 as ASPCAP.
The mean iron abundance in BACCHUS (0.87 0.02) is approximately 0.1 dex lower than ASPCAP (0.78 0.005), yet still aligns with the metal-poor side of the measurements when comparing with the old references. ASPCAP also presents a lower metallicity dispersion than BACCHUS. One of the most notable differences between the ASPCAP and BACCHUS abundances lies in Carbon. Although both methods show the same anti-correlation in C and N and similar trend, the two most Carbon-depleted stars, identified as 2P stars following the [N/Fe] criteria from Geisler et al. (2021), are 0.3-0.4 dex more depleted in BACCHUS than in ASPCAP. For Nitrogen, while ASPCAP and BACCHUS exhibit similar dispersion, ASPCAP reports a mean nitrogen abundance that is 0.1 dex lower than BACCHUS. Regarding the two mentioned 2P stars, their mean nitrogen values are similar in both methods, but the BACCHUS abundances show significantly higher dispersion.
The -elements Si and Ca show a difference of 0.2 in their overall [/Fe] abundance, with BACCHUS reporting a mean of [/Fe] = 0.31 0.02, while ASPCAP reports a mean [/Fe] = 0.13. Oxygen shows similar mean values between ASPCAP ([O/Fe] = 0.28) and BACCHUS ([O/Fe] = 0.26), but BACCHUS exhibits significantly higher dispersion (0.04 vs 0.18). Since MPs are present in this cluster, an oxygen dispersion is expected, lending more credibility to our individual oxygen abundances. Indeed, a N-O anti-correlation (or C-O correlation) is not observed using ASPCAP values (Figure 6).
Magnesium shows a strong agreement in its mean abundance between the two methods, but BACCHUS again exhibits higher dispersion. Notably, two stars in BACCHUS are more magnesium-depleted than any ASPCAP star. Among these, one also displays the high carbon depletion and enhanced nitrogen mentioned earlier, while the other presents a normal level of carbon but enhanced nitrogen abundance (and thus both considered as 2P). Aluminium and Potassium show excellent agreement in both their mean abundance and dispersion across methods. Both ASPCAP and BACCHUS report the same dispersion for Silicon, but BACCHUS abundances are 0.1 dex higher. Calcium in BACCHUS has a slightly higher abundance, with dispersion also being twice as large as in ASPCAP. Titanium shows similar dispersion in both methods but ASPCAP reports abundances that are 0.05 dex higher. However, the typical errors of our Ti measurements allow to determine an agreement between ASPCAP and BACCHUS for this element.
It should be noted that using our [Fe/H] and [/Fe] abundances to calculate the global metallicity [M/H] via the formula from Salaris, Chieffi, & Straniero (1993), we obtain a [M/H]0.3 dex higher than our [Fe/H]. This is expected since GCs are usually -enhanced. Conversely, the global metallicity reported by ASPCAP is almost identical to their iron metallicity, differing by only 0.01 dex, which does not align with their calculation using the aforementioned formula.